OpenStudy (eva12):

Phy CLASS A CHAMPION ATHLETE CAN PRODUCE ONE HORSEPOWER(746 W) FOR A SHORT PERIOD OF TIME. THE NUMBER OF STEPS 16 CM HIGH STEPS A 70 KG ATHLETE COULD ASCEND IN ONE MINTUE WHILE EXPENDING ONE HORSEPOWER IS?tHE COORRCT ANSWER IS=408. THIS IS WHAT I TRIED I COULD NOT GET 408 W=WG=MGH P=W/T=MGH/T G=9.8M/S P=746 W H=16cm M=70KG 746=W=70KG*9.8*16=10796/746=14.7 HOW DO I GET 408 WHAT DID I GO WRONG????

OpenStudy (anonymous):

1st of all: DON'T WRITE IN CAPSLOCK!!!!!!! no one is going to answer... (i seriously don't know, why I am..) 2nd: You made more than one mistake. First of all, you are using the "equal-sign" (=) in places where 2 things are _not_ equal. Don't do that. Where is the time (1 minute) in your last line? Write down what you did.. step by step - don't just write a bunch of stuff and expect not to confuse something.. 3rd. The corrrect answer would in fact be 407, not 408 ;)

OpenStudy (eva12):

step1 equation w=wg=mgh and equation p=w/t=mgh & g=9.8m/s step 2 70kg*9.8m/s*16cm=10976 by using w=mgh step 3 10976/746=14.7

OpenStudy (anonymous):

Okay ahm - you know how your teacher keeps bugging you about "don't plug in the numbers, till the very end" ??? Listen to him. Please - just write down an expression for the # of steps. You will see, where you went wrong.

OpenStudy (anonymous):

And yes, I know: you don't want to bother writing every step down explicitly, because this is an easy one and it's just a tiny bit of arithmetic. True, but the fact, that you keep messing it up, should tell you, you're not capable of doing it in your head. So why do you insist on rushing? Just write it down slow and clean once - you'll be done in a minute..

OpenStudy (eva12):

oo no i was writing the steps on paper sorry i am taking long=[

OpenStudy (anonymous):

That's no problem - take your time. I was just ahm.. expanding what I wrote before... you're doing the right thing. Didn't want to pressure you or anything, sry ;)

OpenStudy (eva12):

p=w/t=mgh step 1 w/t=mgh/p step 2 w=(mgh/p)*t

OpenStudy (eva12):

did i do it wrong????

OpenStudy (anonymous):

Your first line reads\[P = m\, g \, h\]I'm assuming (cause you didn't clarify), that you are saying P is Power. (Should always be a capital P by the way). Are you sure about this?

OpenStudy (eva12):

nope

OpenStudy (anonymous):

Well ahm.. how did you come up with it?

OpenStudy (eva12):

will i was trying to combine equations to try to solve the answer

OpenStudy (anonymous):

that's not really an explanation, what you did is it? If you aren't going to tell us, what you did - we can't really help you.

OpenStudy (eva12):

thanks for your help but i figured it out were i went wrong thanks=]